OpenStudy (anonymous):

10-3x-4x+2x-12 = -2-5x is this right... 10-3x-4x+2x-12 = -2-5x is this right... @Mathematics

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first you combine like terms.. 10-12=-2 3x-4x=-1x -1x+2x=1x so its -2-x..i think this is it..you can go to wolframalpha.com and put in your equation and check